OpenStudy (anonymous):

An object of mass m has an an accleration of magnitude a, when a force of strength F acts on it (on a frictionless surface). Four other objects on the same surface have the following masses and accelerations: A: Mass = 2m, Acceleration = a/2 B: Mass = m, Acceleration = 2a C: Mass = 2m, Acceleration = a D: Mass = 2m, Acceleration = 2a Which of the options below correctly ranks the strengths of the forces acting on these four objects, from weakest to strongest.

OpenStudy (johnweldon1993):

Hint* \(\large F = ma\)

OpenStudy (anonymous):

So all I need to do is multiple the mass and acceleration for all of them? We haven't had this in class yet so I'm not really sure.

OpenStudy (johnweldon1993):

That is correct...And you will be learning about it a lot! Newtons 2nd Law..F = ma So yes just multiply and simplify...what do you get?

OpenStudy (anonymous):

A: 4 B:2 C:2 D:4 Would that be right?

OpenStudy (anonymous):

Because if that's the case how do you decide which one is bigger when B and C are the same values just like A and D?

OpenStudy (johnweldon1993):

Not quite Ex: A) \(\large F = ma\) \(\large F=2m\times \frac{a}{2} = ma\) B) \(\large F = ma\) \(\large F=m\times 2a = 2ma\) *So that should show you that B would be twice as strong of a force as A right? Continue on from there :)

OpenStudy (anonymous):

Sorry about the super late reply. So it would be A<C<D<B?